# Evaluating Large Language Models in Process Mining: Capabilities, Benchmarks, Evaluation Strategies, and Future Challenges

### v1

@@ 1 Introduction
The three main contributions of this paper are summarized in Fig. 1. First, building upon prior work [4, 5] proposing textual abstractions of process mining artifacts and an experimental evaluation of LLMs’ responses, the essential capabilities that LLMs must have for PM tasks are derived in Section 3.1. The aforementioned capabilities allow us to narrow down the field of LLMs to those that meet these requirements. Next, evaluation benchmarks for selecting suitable LLMs are introduced in Section 3.2, incorporating both process-mining-specific and general criteria such as reasoning, visual understanding, factuality, and trustworthiness. Finally, we suggest automatic, human, and self-assessment methods for evaluating LLMs’ outputs on specific tasks in Section 3.3, aiming to establish a comprehensive PM benchmark and enhance confidence in LLMs’ usage, addressing potential issues like hallucination.
@@ 2.2 Implementation Paradigms of Process Mining on LLMs
  • Code generation: LLMs can be used to create structured queries, like SQL, for advanced PM tasks [13]. Rather than directly asking LLMs for answers, users command LLMs to craft database queries from natural language [18]. These queries are then executed on the databases holding PM information. It is applicable to PM tasks that can be converted into database queries, such as filtering event logs or computing the average duration of process steps. Also, LLMs can be used to generate executable programs that use existing PM libraries to infer insights over the event data [11].
@@
LLMs may require additional knowledge about processes and databases to implement PM tasks, for example, in anomaly detection and crafting accurate database queries. Various strategies are used to equip LLMs with this additional domain knowledge [16]. Fine-tuning extends an LLM’s training with a PM-specific dataset, improving task-specific performance [39]. However, it requires significant resources and constant updates. Alternatively, prompt engineering provides a simpler and adaptable method by crafting prompts to guide LLMs’ outputs, incorporating strategies like few-shot learning [6], chain-of-thought [38], and least-to-most [44], without the extensive resource demands of fine-tuning. Additionally, retrieval augmentation techniques leverage external documents to enhance understanding and performance on PM tasks, offering a cost-effective solution for adapting LLMs to specialized analyses.
@@ 3.1 LLMs Capabilities Needed for Process Mining
  • Coding (Text-to-SQL) Capabilities: With the context window limit preventing full event log inclusion in prompts, generating scripts and database queries is crucial for analyzing event data. As discussed in Section 2.2, text-to-SQL assists in filtering and analyzing event data. Key requirements for text-to-SQL in PM include understanding database schemas, performing complex joins, using database-specific operators (e.g., for calculating date differences), and translating PM concepts into queries. Other database languages like Cypher have been explored [30]. Overall, modern LLMs offer excellent coding capabilities [4, 18, 30].
@@ 3.2 Relevant LLMs Benchmarks
In Table 1, we link PM tasks with their implementation paradigms and benchmarks. Below, we iterate over the various tasks described in Table 1.
  • For process description tasks, possessing domain knowledge over the underlying process is key to providing a proper description of the process. For example, in a hospital log, technical terms such as “ER”, “Sepsis”, “ECMO” need to be known by the LLM for a proper understanding of the context.
  • Process modeling involves creating declarative or procedural models from text, and could, for instance, adopt SQL for declarative constraints and BPMN XML for procedural models. It is vital that LLMs can propose multiple hypotheses of process models to an analyst.
  • Anomaly detection and root cause analysis may require domain knowledge of the underlying business process and might be implemented considering the sequence of activities happening in the process, or it may require the LLMs to identify the combinations of event attributes and their values leading to conformance issues.
  • Fairness demands identifying biases in processes, such as in renting, by analyzing event attributes and values for bias detection. It requires the identification of the combinations of event attributes and their values that may be subject to bias. Therefore, being capable of generating valid hypotheses is important.
  • Explaining and interpreting visualizations do not only require understanding a picture (i.e., extracting relevant features) and the annexed textual prompt but also providing contextual insights. For example, knowing how to interpret the crossing of lines of the performance spectrum visualization [17] is required to provide insights into the performance of the underlying process (blocks/batching/bottlenecks).
  • Within process improvement, we can get proposals as text or new declarative/procedural constraints that should be added to the existing process model. Code generation capabilities can be used for this purpose, for instance. The proposition of new constraints requires LLMs to understand the current limitations of the process (performance and biases) and to make hypotheses on how these limitations could be addressed.
@@
Direct Provision
Code Generation
Hypotheses Generation
Traditional
Domain Knowledge
Visual Prompts
Text-to-SQL
Fairness
Hypotheses Generation
Process Description
X
X
X
Process Modeling
X
X
X
X
X
X
X
Anomaly Detection
X
X
X
X
X
X
Root Cause Analysis
X
X
X
X
X
X
Ensuring Fairness
X
X
X
X
X
X
X
Expl. and Interpreting
X
X
X
@@
Process Improvement
X
X
X
X
X
X
X
X
While general-purpose benchmarks are already developed and are easily accessible, they are not entirely suited for the task of PM-on-LLMs. In particular, visual capabilities (explaining and interpreting PM visualizations) and autonomous hypotheses generation require more PM-specific benchmarks. However, little research exists on PM-specific benchmarks.
In [30], a taxonomy of questions to evaluate text-to-SQL and text-to-Cypher is proposed and evaluated on different LLMs. In [4, 5], a set of questions to evaluate the insights directly provided by LLMs is proposed. Moreover, [4] benchmarks LLMs on more advanced multi-prompt and hypotheses generation tasks. In [26], a brief taxonomy of object-centric PM questions is proposed and tested on two LLMs.
@@ 3.3 How to Evaluate LLMs Outputs
  • LLMs Self-Evaluation is useful to mitigate the hallucination problem [28]. Scientific literature suggests several self-evaluation strategies. The ”chain-of-thought” approach necessitates LLMs to detail their reasoning, ensuring explanations for identified causes or anomalies are comprehensive, leading to increased response quality [38]. Another method, the confidence score strategy, involves LLMs assigning confidence levels to each insight, not only aiding users in discerning more reliable insights but also improving the overall insight quality, as LLMs tend to discard low-confidence outputs [32]. Ensembling strategies, which merge outputs from multiple LLMs sessions, also enhance output quality and can be implemented through majority voting or by considering confidence scores from different sessions [21]. Finally, self-reflection strategies are employed, where an LLM’s output is fed into another session of the same or a different LLM for quality evaluation, aiming to identify potential hallucinations in the initial output [29]. For example, in anomaly detection, we can use LLMs to identify semantic anomalies with a confidence score to exclude low-confidence findings, either through the same or different LLM sessions (self-reflection). Additionally, applying the anomaly detection prompt across multiple LLMs sessions (ensembling) yields complementary anomalies, enhancing detection reliability.
4 Research Challenges
Given the results reviewed in this paper, this section aims to highlight some future challenges of evaluating PM-on-LLMs tasks. Particularly, we propose in Section 4.1 the creation of an extensive PM-specific benchmark. In Section 4.2, we identify some challenges related to evaluating LLMs specifically trained on PM information.
4.1 Creation of Extensive PM-specific Benchmarks
In Section 3.2, we considered different general-purpose LLM benchmarks and reported a limited number of PM-specific benchmarks. We also observed that some capabilities, including reasoning on visual prompts and hypotheses generation, are not covered satisfactorily by existing benchmarks. The different evaluation criteria introduced in Section 3.3 help to define the foundations of extensive PM-specific benchmarks. A PM-specific LLM benchmark could cover all the aforementioned strategies in various implementation paradigms, including the direct provision of insights, text-to-SQL, and hypothesis generation (Section 2.2). For the direct provision of insights, annotating each task with anticipated outputs enables the measurement of precision and recall. In the case of text-to-SQL tasks, annotating tasks with expected outcomes facilitates the assessment of the overall correctness of the output beyond just formal accuracy. Regarding hypothesis generation, establishing a thorough set of hypotheses is crucial for evaluating the precision and recall of hypotheses produced by LLMs.
4.2 Tailoring LLMs on PM (and Annexed Benchmarking Challenges)
The currently available LLMs are trained on general-purpose datasets. We argue that the quality of PM-on-LLMs could increase if LLMs trained on PM-specific information were proposed. Some research work is going in this direction. For example, in [16], Large Process Models (LPMs) are proposed using a fine-tuning strategy. In [3], the project ProcessGPT is proposed, which specifically aims to train an LLM on PM knowledge.
However, LLMs tailored for PM tasks could also be difficult to evaluate. A key consideration in benchmarking LLMs for PM is the novelty of the input data, ensuring models are tested on unfamiliar data. Using well-known datasets, like BPI challenge logs, may not accurately test the analytical capabilities of the model because of existing natural language discussions. Therefore, there is a need for research into dynamic, possibly simulated benchmarks that reflect PM’s requirements. These benchmarks should create new scenarios and detect anomalies to better evaluate LLMs’ hypothesis generation and performance in new situations, overcoming the issues of LLMs overfitted on PM.

### v2

# Evaluating Large Language Models in Process Mining: Capabilities, Benchmarks, and Evaluation Strategies

@@ 1 Introduction
The three main contributions of this paper are following. First, building upon prior work [3, 4] proposing textual abstractions of process mining artifacts and an experimental evaluation of LLMs’ responses, the essential capabilities that LLMs must have for PM tasks are derived in Section 3.1. The aforementioned capabilities allow us to narrow down the field of LLMs to those that meet these requirements. Next, evaluation benchmarks for selecting suitable LLMs are introduced in Section 3.2, incorporating both process-mining-specific and general criteria such as reasoning, visual understanding, factuality, and trustworthiness. Finally, we suggest automatic, human, and self-assessment methods for evaluating LLMs’ outputs on specific tasks in Section 3.3, aiming to establish a comprehensive PM benchmark and enhance confidence in LLMs’ usage, addressing potential issues like hallucination.
@@ 2.2 Implementation Paradigms of Process Mining on LLMs
  • Code generation: LLMs can be used to create structured queries, like SQL, for advanced PM tasks [11]. Rather than directly asking LLMs for answers, users command LLMs to craft database queries from natural language. These queries are then executed on the databases holding PM information. It is applicable to PM tasks that can be converted into database queries, such as filtering event logs or computing the average duration of process steps. Also, LLMs can be used to generate executable programs that use existing PM libraries to infer insights over the event data [9].
@@
LLMs may require additional knowledge about processes and databases to implement PM tasks, for example, in anomaly detection and crafting accurate database queries. Some strategies are used to equip LLMs with this additional domain knowledge [14], including fine-tuning and prompt engineering.
@@ 3.1 LLMs Capabilities Needed for Process Mining
  • Coding (Text-to-SQL) Capabilities: With the context window limit preventing full event log inclusion in prompts, generating scripts and database queries is crucial for analyzing event data. As discussed in Section 2.2, text-to-SQL assists in filtering and analyzing event data. Key requirements for text-to-SQL in PM include understanding database schemas, performing complex joins, using database-specific operators (e.g., for calculating date differences), and translating PM concepts into queries. Overall, modern LLMs offer excellent coding capabilities [3].
@@ 3.2 Relevant LLMs Benchmarks
In Table 1, we link process mining (PM) tasks to implementation paradigms and benchmarks. We discuss these tasks:
  • Process description requires understanding technical terms relevant to the domain, crucial for accurately describing processes.
  • Process modeling involves generating models from text, using SQL for declarative and BPMN XML for procedural models. LLMs should offer various model hypotheses.
  • Anomaly detection and root cause analysis need domain knowledge to analyze process sequences or identify event attribute combinations causing issues.
  • Fairness involves detecting biases by analyzing event attributes and values, necessitating hypothesis generation by LLMs.
  • Explaining and interpreting visualizations requires extracting features from images and texts, offering contextual insights, like interpreting performance spectrum visualization [15].
  • Process improvement entails suggesting text proposals or new constraints to enhance current models, leveraging code generation capabilities and understanding process limitations.
@@
Direct Provision Code Generation Hypotheses Generation Traditional Domain Knowledge Visual Prompts Text-to-SQL Fairness Hypotheses Generation
Process Description X X X
Process Modeling X X X X X X X
Anomaly Detection X X X X X X
Root Cause Analysis X X X X X X
Ensuring Fairness X X X X X X X
Expl. and Interpreting X X X
@@
Process Improvement X X X X X X X X
While general-purpose benchmarks are already developed and are easily accessible, they are not entirely suited for the task of PM-on-LLMs. In particular, visual capabilities (explaining and interpreting PM visualizations) and autonomous hypotheses generation require more PM-specific benchmarks. However, little research exists on PM-specific benchmarks [3, 4].
@@ 3.3 How to Evaluate LLMs Outputs
  • Self-evaluation in large language models (LLMs) addresses the hallucination issue, as highlighted by [24]. Various strategies include the chain-of-thought approach, where LLMs elaborate on their reasoning processes, improving explanation quality [33]. Confidence scores allow LLMs to rate their insights’ reliability, encouraging higher quality by discarding less certain outputs [27]. Ensembling, combining results from several LLM sessions through majority voting or confidence evaluation, boosts output accuracy [18]. Self-reflection, where an LLM reviews its output or that of another model, helps catch inaccuracies [25]. In practices like anomaly detection, leveraging confidence scores to filter out uncertain anomalies and ensembling for corroborating detection across sessions enhance reliability.